Complete the ordered pairs so that each is a solution for the given equation.
3x+4y=12 (0, ), ( ,3/4), ( ,0), (5/3, )

Answer by pvaka(23)

3x+4y=12 (0,3), (3,3/4), (4,0), (5/3,7/4 )
3(0)+4y=12
0+4y=12
4y/4=12/4
y=3
3x+4(3/4)=12
3x+12/4=12
3x+3-3=12-3
3x=9
3x/3-9/3
x=3
3x+4(0)=12
3x+0=12
3x=12
3x/3-12/3
x=4
3(5/3)+4y=12
15/3+4y=12
5+4y-5=12-5
4y=7
4y/4=7/4
y=7/4